# FedShard: Federated Unlearning with Efficiency Fairness and Performance Fairness

### v1

@@ 1 Introduction
Federated learning (FL) is a prominent paradigm for collaborative machine learning on distributed data, wherein decentralized clients train a shared global model under the coordination of a central server (McMahan et al. 2017). To protect the “right to be forgotten” (RTBF) for clients who leave the FL system, federated unlearning (FU) has emerged as an essential method for removing the contributions of leaving data from the global model (Voigt and Von dem Bussche 2017; Harding et al. 2019; Truex et al. 2019; Mothukuri et al. 2021). However, prior work in FU mainly focuses on the efficiency and effectiveness of the unlearning process (Nguyen et al. 2022; Zhang et al. 2023a). We argue that the fairness of federated unlearning—with respect to both computational costs and performance impacts—is a critical yet overlooked aspect, especially for system sustainability and security.
As illustrated in Figure 1, literature usually overlooked two types of issues, performance fairness (P.F.) and efficiency fairness (E.F.), which are crucial and practically important. P.F. aims to achieve a fair performance degradation caused by unlearning among clients, guaranteeing both remaining clients’ rights to good model performance, and leaving clients’ RTBF. Without P.F., unlearning leaving clients may cause an excessive accuracy decrease on specific data, thereby encouraging other remaining clients (especially those with similar data) to leave. Such cascaded leaving can severely compromise the sustainability of an FL system. Moreover, unfair unlearning can be further exploited for data poisoning attacks (Tolpegin et al. 2020; Nuding and Mayer 2022). By issuing unlearning requests intentionally, malicious clients can launch data poisoning attacks on specific clients with similar data and compromise the security of the FL system. E.F. demands a fair unlearning cost for every client when it leaves. Unlearning methods without E.F. may harm the system sustainability by discouraging clients from participating in the system, as they are concerned about unfairly large leaving costs.
@@
Current FU methods are unsatisfactory in terms of performance fairness and efficiency fairness. Specifically, related work can be classified into two categories: calibration-based approximate unlearning (Li et al. 2023; Lin et al. 2024; Gao et al. 2024; Tarun et al. 2024) and retraining-based exact unlearning (Xiong et al. 2023). As illustrated in Figure 2(a), calibration unlearning (such as FedRecover (Zhang et al. 2023b), FedEraser (Liu et al. 2021a) and RapidRetrain (Liu et al. 2022)) often introduces extra performance degradation on remaining clients or insufficient unlearning on leaving clients, which compromises the P.F. Exact retraining (such as FATS (Tao et al. 2024)) can easily introduce significant variance in unlearning efficiency, thereby violating E.F..
@@
Figure 2: Left: An example of the FedShard architecture with a merging rate of 3. Clients are organized in a tree-like structure. When clients request unlearning (indicated in blue), only the shards on the direct path to the root (dark background) require retraining. Right: A conceptual illustration showing that only FedShard simultaneously achieves both performance fairness and efficiency fairness.
@@ 2 Related Work
Parameter calibration methods scale down leaving clients’ gradients with the Hessian matrix (Liu et al. 2022), cached intermediate information (Liu et al. 2021a), or gradient ascent (Tarun et al. 2024). They achieve efficient federated unlearning, while such methods can easily cause extra performance degradation (Zhang et al. 2023b) on remaining clients and insufficient unlearning on leaving clients, violating P.F. Our proposed method will avoid such issues and achieve performance fairness while maintaining efficiency.
Exact retraining methods achieve quite effective unlearning, e.g., FATS (Tao et al. 2024) and Exact-fun (Xiong et al. 2023). They utilize cached information to alleviate the heavy overhead caused by vanilla retraining. Unfortunately, these methods usually lead to disparities in unlearning costs, compromising efficiency fairness. For example, Tao et al. allow clients to join at different rounds for fast federated unlearning. When the later-joined client is leaving, earlier rounds without its participation can be bypassed from unlearning, thereby achieving efficient unlearning. However, earlier-joined clients will burden extra overheads, leading to unfair unlearning costs. Our FedShard ensures that every client has a closer unlearning cost related only to its own data contribution, thereby ensuring a fair unlearning cost while maintaining efficient retraining.
@@ 4.1 Effectiveness of Unlearning
We hold the following proposition that is widely accepted and leveraged in the exact unlearning literature (Bourtoule et al. 2021; Tao et al. 2024; Xiong et al. 2023):
@@ 5.3 Fairness Properties
Our proposed metrics also satisfy several desirable properties established in economic and computer science literature, including envy-freeness (FP.1), Pareto optimality (FP.2), fairness axiom (FP.3), and reduction (FP.4) (Li and Xue 2013a; Procaccia 2013; Feldman and Kirman 1974; Luc 2008; Lan et al. 2010; Li and Xue 2013b). FU algorithms minimizing $M_{p}$ satisfy FP.1, FP.2, and FP.3; FU algorithms minimizing $M_{e}$ satisfy FP.3 and FP.4. These properties ensure the robustness and theoretical grounding of our metrics. We provide formal definitions and proofs that FedShard satisfies these properties in Appendix F.
@@ 6.3 Effectiveness and Performance Fairness
P.F. evaluation via DPA Attack experiment. If an FU algorithm is not performance-fair (insufficient unlearning or over-unlearning), it can be easily attacked by data poisoning attacks (DPA) (Tolpegin et al. 2020; Nuding and Mayer 2022; Gupta et al. 2023; Wang and Chaudhuri 2018) as we explain in Appendix G.5. As shown in Table 2, the approximate unlearning methods RR and FE fail to defend against DPA and exhibit high $M_{p}$ scores, consistent with the performance unfairness observed in the cascaded leaving experiment. In contrast, the retraining methods FS, FT, and FA successfully defend against DPA while achieving low $M_{p}$ scores, indicating that FS performs well in terms of P.F..
@@ A.1 Fairness in Federated Unlearning
Existing literature focused on data fairness and model fairness, which pay attention to whether specific groups of people suffer from racial discrimination (Oesterling et al. 2024; Shao et al. 2024; Le Quy et al. 2022; Oneto and Chiappa 2020; Pessach and Shmueli 2022) or whether data providers with different data qualities receive the same rewards (Su et al. 2024). These notions of fairness are orthogonal to the concepts we introduce in this paper. Our work is the first to formalize and address fairness within the unlearning mechanism itself, focusing on the equitable distribution of performance impacts (P.F.) and computational burdens (E.F.) across the client population.
@@ A.2 Sharded Unlearning
The concept of sharding, or partitioning data to facilitate unlearning, has been explored previously in different contexts. SISA (Sharded, Isolated, Sliced, and Aggregated) (Bourtoule et al. 2021) introduced a sharded architecture for centralized machine learning, where data is divided into mutually exclusive shards, and retraining is confined to only the shard containing the data to be forgotten. However, its reliance on a central server with direct data access makes it incompatible with the privacy-preserving constraints of FL.
More recently, ShardingEraser (Lin et al. 2024) applied sharding to FL and utilized sharding to accelerate an approximate unlearning method (FedEraser). Consequently, it inherits the fundamental limitations of approximate approaches and fails to guarantee performance fairness. In contrast, FedShard is the first framework to leverage a novel hierarchical sharding structure to achieve exact unlearning in the FL setting, while co-optimizing for both performance and efficiency fairness.
@@ Appendix C Contribution Factor
In Equation (1), we utilize the parameter similarity to measure the contribution factor $\alpha_{c}$ for every client, which is widely used in FL literature for adaptive aggregation (Wu and Wang 2021) and personalization (Huang et al. 2021). With the cosine value of the model parameters, we capture adequate information about the data contribution of each client while having a small computational cost.
The contribution factor can also be measured by the loss similarity. The loss similarity utilizes the loss value of the model to measure the contribution factor of each client. For instance, we calculate the loss values on every client’s local model. Lower loss value indicates higher contribution. This rough but fast method has already been used in personalization (Zhang et al. 2020). As a metric for the contribution factor, while this can be marginally faster to compute than vector similarity in then existing design, it is a scalar metric and thus less expressive. Two clients could reduce the loss by a similar amount but through conflicting updates; this critical nuance is captured by parameter similarity but lost when using only the loss value.
The Shapley value is an axiomatic method from cooperative game theory that provides a theoretically sound and provably fair attribution of contribution to each client (Winter 2002). This is also leveraged for personalization works to determine the data relevance (Wu et al. 2024). The Shapley value is almost the theoretical “gold standard” for fair contribution attribution. However, its exact computation is NP-hard, and even state-of-the-art approximations are prohibitively expensive in large-scale FL systems. Its computational demands make it impractical for a dynamic and efficient framework like FedShard.
The information entropy utilizes information-theoretic measures to quantify a client’s contribution, which has been proven to be practical in adaptive aggregation works (Zheng et al. 2023). This metric offers a different and valuable perspective on contribution, particularly by identifying clients that provide unique information. However, as a contribution factor, its computation is generally much more intensive than parameter similarity.
@@ D.1 Federated Learning Algorithm of FedShard
where $\mathcal{L}_{c}(\cdot)$ is the loss function for client $c$. To achieve this, each shard starts federated training among its clients $\mathcal{C}_{s}^{p}$ for $T_{s}^{p}$ rounds, with the initial global model $\theta^{p,0}_{s}$. In this work, we use FedAvg (McMahan et al. 2017) as the federated training algorithm, which is widely used in FL.
@@ D.3 Training Round Allocation Algorithm $\mathcal{A}_{2}$
Variance-based round allocation: With the range determined, we adjust the training rounds for each shard in the subsequent stages. Existing literature (Wu and Wang 2021; Tu et al. 2022; Zheng et al. 2023; Liu et al. 2021b) demonstrates that adaptive assignment of training resources (e.g., learning rates, aggregation weights, etc.) to clients who contribute more is crucial for accelerating convergence in federated learning and unlearning. Therefore, for the remaining stages, we determine the training rounds $T_{s}^{p}$ for each shard $s$ in the same stage $p$ based on the following formula:
@@ E.1 Effectiveness of FedShard unlearning
In this section, we prove that FedShard unlearning, which only trains the one shard containing the leaving client every stage, is as effective as retraining from scratch. Proposition 1 is well-established in the literature (Bourtoule et al. 2021; Tao et al. 2024; Xiong et al. 2023; Lin et al. 2024), stating that for any isolated subsystem in machine learning, if it is unaffected by departing clients, we can directly use cached updates without retraining. We define that $Y$ is affected by $X$ if and only if changes in $X$ lead to changes in $Y$.
@@ G.1 Evaluation Settings
Datasets: We evaluate FedShard on well-known datasets including MNIST, Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), CIFAR-10, CIFAR-100, EMNIST (Cohen et al. 2017), SVHN (Netzer et al. 2011), Purchase, and Adult.
Non-IID: To simulate the non-IID data distribution in FL scenarios, we apply the Label-based Dirichlet Partition (LDA) (Hsu, Qi, and Brown 2019), a method that divides data into multiple subsets based on label distributions using the principles of the Dirichlet process with $\rho$ as the hyperparameter. A lower $\rho$ indicates a higher non-IID degree. In this experiment, we take $\rho=0.1$ to simulate a high non-IID case, $\rho=0.5$ for a moderate non-IID case, and $\rho=0.9$ for an IID case.
@@ G.5 MIA Attacks and Cascaded Leaving
Membership Inference Attacks (MIA) is a common method used in related works to empirically show the effectiveness of federated unlearning (Zhang et al. 2023b; Liu et al. 2022; Tao et al. 2024; Liu et al. 2021a; Lin et al. 2024; Bourtoule et al. 2021; Yan et al. 2022; Dukler et al. 2023; Chen et al. 2024; Hoang et al. 2024; Tu, Hu, and Ma 2024). In MIA, the attacker queries the model with the target sample and the shadow dataset. After that, the attacker compares confidence scores/behavior on member and non-member data to determine if a specific data sample was used to train a model. If one federated unlearning method is effective, it is necessary for it to defend against MIA attacks. Usually, the MIA attacks are evaluated by the F1 score, which is the harmonic mean of precision and recall. The higher the F1 score, the more successful the MIA attacks are.
Cascaded leaving in FL: Literature (Ding et al. 2023; Ding, Wei, and Berry 2024; Zhang, Wei, and Berry 2021; Zhan et al. 2020) shows that strategic clients may leave the FL systems after some clients unlearn their data, which is called cascaded leaving. When one client $c_{u}$ is leaving though federated unlearning, the payoffs of another client $c_{r}$ is $W_{r}^{k}$$\Delta\mathcal{Y}-Z+\lambda_{r}$ if $c_{r}$ keeps remaining, and $W_{r}^{l}$$\mathcal{Y}^{*}+\lambda_{r}$ if $c_{r}$ leaves together, where $\Delta\mathcal{Y}$ denotes the performance degradation, $Z$ denotes unlearning cost, $\lambda_{r}$ represents other payoffs irrelevant to leaving or remaining decisions such as training costs. We do not need to measure the value of $\lambda_{r}$ when just comparing the difference of payoffs between leaving and remaining, i.e., $W_{r}^{k}$ and $W_{r}^{l}$. Client $c_{r}$ will also leave the system when $W_{r}^{k}<W_{r}^{l}$, i.e., $\Delta Y>Y^{*}-Z$. If a federated unlearning algorithm is unfair, the large model performance degradation will intensify cascaded leaving.

### v2

@@ 1 Introduction
Federated learning (FL) is a prominent paradigm for collaborative machine learning on distributed data, wherein decentralized clients train a shared global model under the coordination of a central server (fedavg). To protect the “right to be forgotten” (RTBF) for clients who leave the FL system, federated unlearning (FU) has emerged as an essential method for removing the contributions of leaving data from the global model (GDPR; CCPA; privacyconcerns; FLsurvey). However, prior work in FU mainly focuses on the efficiency and effectiveness of the unlearning process (mulsurvey; mulsurvey2). We argue that the fairness of federated unlearning—with respect to both computational costs and performance impacts—is a critical yet overlooked aspect, especially for system sustainability and security.
As illustrated in Figure 1, literature usually overlooked two types of issues, performance fairness (P.F.) and efficiency fairness (E.F.), which are crucial and practically important. P.F. aims to achieve a fair performance degradation caused by unlearning among clients, guaranteeing both remaining clients’ rights to good model performance, and leaving clients’ RTBF. Without P.F., unlearning leaving clients may cause an excessive accuracy decrease on specific data, thereby encouraging other remaining clients (especially those with similar data) to leave. Such cascaded leaving can severely compromise the sustainability of an FL system. Moreover, unfair unlearning can be further exploited for data poisoning attacks (dpa1). By deliberate unlearning requests, malicious clients can launch data poisoning attacks on specific clients with similar data and compromise the security of the FL system. E.F. demands a fair unlearning cost for every client when it leaves. Unlearning methods without E.F. may harm the system sustainability by discouraging clients from participating in the system, as they are concerned about unfairly large leaving costs.
@@
Current FU methods are unsatisfactory in terms of performance fairness and efficiency fairness. Specifically, related work can be classified into two categories: calibration-based approximate unlearning (FedAF; ShardingE; verifi; FYEFUL) and retraining-based exact unlearning (exactfun). As illustrated in Figure 2(a), calibration unlearning (such as FedRecover (fedrec), FedEraser (federaser) and RapidRetrain (RR)) often introduces extra performance degradation on remaining clients or insufficient unlearning on leaving clients, which compromises the P.F. Exact retraining (such as FATS (FATS)) can easily introduce significant variance in unlearning efficiency, thereby violating E.F..
@@
Figure 2: Left: A conceptual illustration showing that only FedShard simultaneously achieves both performance fairness and efficiency fairness. Right: An example of the FedShard architecture with a merging rate of 3. Clients are organized in a tree-like structure. When clients request unlearning (indicated in blue), only the shards on the direct path to the root (dark background) require retraining.
@@ 2 Related Work
Parameter calibration methods scale down leaving clients’ gradients with the Hessian matrix (RR), cached intermediate information (federaser), or gradient ascent (FYEFUL). They achieve efficient federated unlearning, while such methods can easily cause extra performance degradation (fedrec) on remaining clients and insufficient unlearning on leaving clients, violating P.F. Our proposed method will avoid such issues and achieve performance fairness while maintaining efficiency.
Exact retraining methods achieve quite effective unlearning, e.g., FATS (FATS) and Exact-fun (exactfun). They utilize cached information to alleviate the heavy overhead caused by vanilla retraining. Unfortunately, these methods usually lead to disparities in unlearning costs, compromising efficiency fairness. For example, FATS allow clients to join at different rounds for fast federated unlearning. When the later-joined client is leaving, earlier rounds without its participation can be bypassed from unlearning, thereby achieving efficient unlearning. However, earlier-joined clients will burden extra overheads, leading to unfair unlearning costs. Our FedShard ensures that every client has a closer unlearning cost related only to its own data contribution, thereby ensuring a fair unlearning cost while maintaining efficient retraining.
@@ 4.1 Effectiveness of Unlearning
We hold the following proposition that is widely accepted and leveraged in the exact unlearning literature (SISA; FATS; exactfun):
@@ 5.3 Fairness Properties
Our proposed metrics also satisfy several desirable properties established in economic and computer science literature, including envy-freeness (FP.1), Pareto optimality (FP.2), fairness axiom (FP.3), and reduction (FP.4) (fairpro1; fairpro2; envyfree1; pareto; axiom; fair-eco). FU algorithms minimizing $M_{p}$ satisfy FP.1, FP.2, and FP.3; FU algorithms minimizing $M_{e}$ satisfy FP.3 and FP.4. These properties ensure the robustness and theoretical grounding of our metrics. We provide formal definitions and proofs that FedShard satisfies these properties in Appendix F.
@@ 6.3 Effectiveness and Performance Fairness
P.F. evaluation via DPA Attack experiment. If an FU algorithm is not performance-fair (insufficient unlearning or over-unlearning), it can be easily attacked by data poisoning attacks (DPA) (dpa1; dpa2; dpa3; dpa4) as we explain in Appendix G.5. As shown in Table 2, the approximate unlearning methods RR and FE fail to defend against DPA and exhibit high $M_{p}$ scores, consistent with the performance unfairness observed in the cascaded leaving experiment. In contrast, the retraining methods FS, FT, and FA successfully defend against DPA while achieving low $M_{p}$ scores, indicating that FS performs well in terms of P.F..
@@ 7 Conclusion
8 Acknowledgments
This work is supported by National Natural Science Foundation of China under Grant 62502412. This research is also partially supported by National Key R&D Program of China under Grant 2024YFE0200500 and National Natural Science Foundation of China under Grants 62202427 and 62572433.
@@ A.1 Fairness in Federated Unlearning
Existing literature focused on data fairness and model fairness, which pay attention to whether specific groups of people suffer from racial discrimination (ful; fairFUL; dfairML2; dfairML3; fairMU) or whether data providers with different data qualities receive the same rewards (f2ul). These notions of fairness are orthogonal to the concepts we introduce in this paper. Our work is the first to formalize and address fairness within the unlearning mechanism itself, focusing on the equitable distribution of performance impacts (P.F.) and computational burdens (E.F.) across the client population.
@@ A.2 Sharded Unlearning
The concept of sharding, or partitioning data to facilitate unlearning, has been explored previously in different contexts. SISA (Sharded, Isolated, Sliced, and Aggregated) (SISA) introduced a sharded architecture for centralized machine learning, where data is divided into mutually exclusive shards, and retraining is confined to only the shard containing the data to be forgotten. However, its reliance on a central server with direct data access makes it incompatible with the privacy-preserving constraints of FL.
More recently, ShardingEraser (ShardingE) applied sharding to FL and utilized sharding to accelerate an approximate unlearning method (FedEraser). Consequently, it inherits the fundamental limitations of approximate approaches and fails to guarantee performance fairness. In contrast, FedShard is the first framework to leverage a novel hierarchical sharding structure to achieve exact unlearning in the FL setting, while co-optimizing for both performance and efficiency fairness.
A.3 Other Related Work
Hierarchical federated learning introduce multiple aggregation layers according to the physical network structure to reduce communication costs and improve scalability (hfl; hfl1). FedShard, however, employs a logical hierarchical sharding structure to facilitate efficient and fair unlearning, which is orthogonal to the physical hierarchy in HFL. Moreover, original HFL’s network/topology-based structure damages the E.F. required by FedShard, i.e., leaving clients on the critical path inevitably incurs high costs.
Clustered federated learning handle non-IID data by partitioning clients into clusters based on similarity and then training separate models for each cluster (cfl1). However, FedShard’s target is different in clustered federated learning scenarios, which lacks a unified global model for exact unlearning.
There are also machine unlearning literature on tree-based structures (DynFrs). However, they focus on machine unlearning for random forest, where the tree-like structure serves as the model structure rather than FedShard, where a data partitioning strategy for federated unlearning.
@@ Appendix C Contribution Factor
In Equation (1), we utilize the parameter similarity to measure the contribution factor $\alpha_{c}$ for every client, which is widely used in FL literature for adaptive aggregation (fedadp) and personalization (FedAMP). With the cosine value of the model parameters, we capture adequate information about the data contribution of each client while having a small computational cost.
The contribution factor can also be measured by the loss similarity. The loss similarity utilizes the loss value of the model to measure the contribution factor of each client. For instance, we calculate the loss values on every client’s local model. Lower loss value indicates higher contribution. This rough but fast method has already been used in personalization (FedFomo). As a metric for the contribution factor, while this can be marginally faster to compute than vector similarity in then existing design, it is a scalar metric and thus less expressive. Two clients could reduce the loss by a similar amount but through conflicting updates; this critical nuance is captured by parameter similarity but lost when using only the loss value.
The Shapley value is an axiomatic method from cooperative game theory that provides a theoretically sound and provably fair attribution of contribution to each client (shapley). This is also leveraged for personalization works to determine the data relevance (pFedSV). The Shapley value is almost the theoretical “gold standard” for fair contribution attribution. However, its exact computation is NP-hard, and even state-of-the-art approximations are prohibitively expensive in large-scale FL systems. Its computational demands make it impractical for a dynamic and efficient framework like FedShard.
The information entropy utilizes information-theoretic measures to quantify a client’s contribution, which has been proven to be practical in adaptive aggregation works (fedent). This metric offers a different and valuable perspective on contribution, particularly by identifying clients that provide unique information. However, as a contribution factor, its computation is generally much more intensive than parameter similarity.
@@ D.1 Federated Learning Algorithm of FedShard
where $\mathcal{L}_{c}(\cdot)$ is the loss function for client $c$. To achieve this, each shard starts federated training among its clients $\mathcal{C}_{s}^{p}$ for $T_{s}^{p}$ rounds, with the initial global model $\theta^{p,0}_{s}$. In this work, we use FedAvg (fedavg) as the federated training algorithm, which is widely used in FL.
@@ D.3 Training Round Allocation Algorithm $\mathcal{A}_{2}$
Variance-based round allocation: With the range determined, we adjust the training rounds for each shard in the subsequent stages. Existing literature (fedadp; fedent; asFL2) demonstrates that adaptive assignment of training resources (e.g., learning rates, aggregation weights, etc.) to clients who contribute more is crucial for accelerating convergence in federated learning and unlearning. Therefore, for the remaining stages, we determine the training rounds $T_{s}^{p}$ for each shard $s$ in the same stage $p$ based on the following formula:
@@ E.1 Effectiveness of FedShard unlearning
In this section, we prove that FedShard unlearning, which only trains the one shard containing the leaving client every stage, is as effective as retraining from scratch. Proposition 1 is well-established in the literature (SISA; FATS; exactfun; ShardingE), stating that for any isolated subsystem in machine learning, if it is unaffected by departing clients, we can directly use cached updates without retraining. We define that $Y$ is affected by $X$ if and only if changes in $X$ lead to changes in $Y$.
@@ G.1 Evaluation Settings
Datasets: We evaluate FedShard on well-known datasets including MNIST, Fashion-MNIST (fmnist), CIFAR-10, CIFAR-100, EMNIST (emnist), SVHN (SVHN), Purchase, and Adult.
Non-IID: To simulate the non-IID data distribution in FL scenarios, we apply the Label-based Dirichlet Partition (LDA) (dirichlet), a method that divides data into multiple subsets based on label distributions using the principles of the Dirichlet process with $\rho$ as the hyperparameter. A lower $\rho$ indicates a higher non-IID degree. In this experiment, we take $\rho=0.1$ to simulate a high non-IID case, $\rho=0.5$ for a moderate non-IID case, and $\rho=0.9$ for an IID case.
@@ G.5 MIA Attacks and Cascaded Leaving
Membership Inference Attacks (MIA) is a common method used in related works to empirically show the effectiveness of federated unlearning (fedrec; RR; FATS; federaser; ShardingE; SISA; review1; review2; review3; review4; game). In MIA, the attacker queries the model with the target sample and the shadow dataset. After that, the attacker compares confidence scores/behavior on member and non-member data to determine if a specific data sample was used to train a model. If one federated unlearning method is effective, it is necessary for it to defend against MIA attacks. Usually, the MIA attacks are evaluated by the F1 score, which is the harmonic mean of precision and recall. The higher the F1 score, the more successful the MIA attacks are.
Cascaded leaving in FL: Literature (IncentiveD1; IncentiveD2; IncentiveD3; IncentiveZ; Incentive) shows that strategic clients may leave the FL systems after some clients unlearn their data, which is called cascaded leaving. When one client $c_{u}$ is leaving though federated unlearning, the payoffs of another client $c_{r}$ is $W_{r}^{k}$$\Delta\mathcal{Y}-Z+\lambda_{r}$ if $c_{r}$ keeps remaining, and $W_{r}^{l}$$\mathcal{Y}^{*}+\lambda_{r}$ if $c_{r}$ leaves together, where $\Delta\mathcal{Y}$ denotes the performance degradation, $Z$ denotes unlearning cost, $\lambda_{r}$ represents other payoffs irrelevant to leaving or remaining decisions such as training costs. We do not need to measure the value of $\lambda_{r}$ when just comparing the difference of payoffs between leaving and remaining, i.e., $W_{r}^{k}$ and $W_{r}^{l}$. Client $c_{r}$ will also leave the system when $W_{r}^{k}<W_{r}^{l}$, i.e., $\Delta Y>Y^{*}-Z$. If a federated unlearning algorithm is unfair, the large model performance degradation will intensify cascaded leaving.